# Pólya’s conjecture up to $\epsilon$-loss and quantitative estimates for the remainder of Weyl’s law 00footnotetext: 2020 Mathematics Subject Classification. Primary 35P15; Secondary 35P20, 42B37. Key words and phrases: Dirichlet eigenvalue, Weyl law, Pólya conjecture

### v1

@@
where $\Lambda(\epsilon,\Omega)$ is given explicitly. This reduces the $\epsilon$-loss version of Pólya’s conjecture to a computational problem. This estimate is based on a uniform estimate on the remainder of the Weyl law on Lipschitz domains, which appears to be the first quantitative estimate for the remainder of Weyl’s law since Weyl’s seminal work in the year 1911. We also provide in all dimensions $n\geq 2$, a class of domains that may even have rather irregular shapes or boundaries but satisfy Pólya’s conjecture.
@@ 1.1 Weyl’s law and Pólya’s conjecture
The Weyl law is also important in the longstanding question “can one hear about the shape of a cavity”, see Lord Rayleigh [27] and Kac [21]. Moreover, the Weyl law is even related to the number theory, Prime number theorem, Riemann hypothesis, via Connes’ trace formula, see Connes [4], Fathizadeh-Khalkhali [11] and the beamer by Khalkhali [22].
@@
Subsequent to Weyl’s seminal work, Courant [5], Courant-Hilbert [6], Hörmander [17, 18], Seeley [32, 33] and many others (see Ivrii [20] for comprehensive description on related history) made contributions to improve the estimate on the remainder. Up to present, the best estimate is as $O(\lambda^{\frac{n-1}{2}})$ on piecewise smooth domains by Seeley [32, 33], see also Hörmander [18].
@@
However, none of the above mentioned results gave quantitative estimate for the remainder $R_{\Omega}(\lambda)$ (that is, these results states that certain asymptotic estimates valid for large $\lambda$’s without quantified the sizes of such $\lambda$’s), and consequently they do not give any quantitative estimate on eigenvalues. From both mathematical and physical point of view, it is rather natural to ask the following question.
@@
From the mathematical point of view, uniform and quantitative estimates are always more interesting. In fact, in 1954, Pólya [29] conjectured that, it should hold on arbitrary bounded open domain that
@@
The Pólya conjecture has been wide open since then. Other than tiling domains, Laptev [24] verified Pólya’s conjecture on product domains $\Omega_{1}\times\Omega_{2}\subset{\mathbb{R}}^{n_{1}+n_{2}}$, $n_{1}\geq 2$, $n_{2}\geq 1$, provided Pólya’s conjecture holds on $\Omega_{1}$. Filonov et al. [7] and [8] verified, via a method from number theory, Pólya’s conjecture on balls for all $n\geq 2$, and on annuli for $n=2$. For general domains $\Omega$, it was known that Pólya’s conjecture holds for the first two Dirichlet eigenvalues see [16] for the Rayleigh-Faber-Krahn inequality ($\lambda_{1}(\Omega)$) and the Krahn-Szegö inequality ($\lambda_{2}(\Omega)$). See also [16] for the first nontrivial Neumann eigenvalue and recent exciting developments [3, 13] for the second nontrivial Neumann eigenvalue.
@@
Our main purpose of this paper is to prove Pólya’s conjecture up to an $\epsilon$-loss, and to give a solution to Question 1.1, on Lipschitz domains.
1.2 Main results
Let us first give a definition for Pólya’s conjecture up to an $\epsilon$-loss.
@@ 1.2 Main results
Noting that $2\leq\left(\frac{n+2}{n}\right)^{n/2}\to e$ as $2\leq n\to\infty$, the $\epsilon$-loss version of Pólya’s conjecture improves substantially (1.5) and its improvements on lower order terms, see [2, 12, 23, 28].
@@
Our first main result is as following.
@@
Obviously, for a given $\epsilon\in(0,1)$, and a concrete Lipschitz domain $\Omega$, $\Lambda(\epsilon,\Omega)$ is uniquely determined by the equation (1.7). Hence, to show the $\epsilon$-loss version of Pólya’s conjecture, Theorem 1.3 reduces the problem to a computational problem for eigenvalues $\lambda_{k}(\Omega)<\Lambda(\epsilon,\Omega)$.
@@
We have the following uniform estimate on Lipschitz domains, which gives a solution to Question 1.1. Moreover, our uniform estimate appears to be the first quantitative estimate for the remainder of Weyl’s law since Weyl’s seminal work [34] in the year 1911, see also Courant [5], Courant-Hilbert [6], Hörmander [18], Ivrii [19], Seeley [32, 33] etc for asymptotic results. In what follows, for a bounded open domain $\Omega$, we denote by $r_{in}=r_{in}(\Omega)$ the radius of the largest ball contained in $\Omega$.
@@
$k$ $\geq\frac{\omega(n)|\Omega|}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2}-\frac{C_{Lip}(\Omega)\omega(n)|\partial\Omega|}{(2\pi)^{n}}\lambda_{k}(\Omega)^{\frac{n-1}{2}}\left(\frac{\pi r_{in}(\Omega)}{|I_{min}|}+5n^{2}\pi\log_{2}\frac{10|I_{min}|\sqrt{\lambda_{k}(\Omega)}}{\pi}\right).$
@@
Remark 1.6.
If $\Omega\subset{{{\mathbb{R}}}^{n}}$ is a convex set, then one can take $C_{Lip}(\Omega)$ to be 1 in the above two theorems. In fact, one can have better estimates than (1.7) and (1.8). See Remark (4.2) below.
Obviously, both quantities $\frac{r_{in}(\mathscr{R}\setminus\overline{\Omega})}{|I_{min}|}$ and $\frac{r_{in}(\Omega)}{|I_{min}|}$ are not larger than $1/2$, and therefore can be replaced by $1/2$. Moreover, note that if $|\partial(\mathscr{R}\setminus\overline{\Omega})|$ is small compared to $|\mathscr{R}_{n-1}|$, we then have that Pólya’s conjecture holds for the first some eigenvalues. Our quantitative estimate is consistent to the asymptotic estimates established by Courant [5] and Courant-Hilbert [6], which is, up to present, best possible in the case of domains having only Lipschitz continuous boundary.
@@
1.3 Pólya’s conjecture on a class of domains
Using the strategy developed above, we can provide in all dimensions $n\geq 2$, a class of domains which may have rather irregular shape or boundary but satisfy Pólya’s conjecture.
@@ 1.3 Pólya’s conjecture on a class of domains
Suppose $\mathscr{R}=\mathscr{R}_{n-1}\times I$ is a rectangle in ${{{\mathbb{R}}}^{n}}$, $n\geq 2$. Let $\{Q_{k}\}_{k\in{\mathbb{N}}}$ be of Admissible class in ${{{\mathbb{R}}}^{n}}$, that is contained in $\mathscr{R}$. If $|\mathscr{R}_{n-1}|>C_{2}(n-1)\mathscr{S}_{Q}$, then the Pólya conjecture holds on $\mathscr{R}\setminus\overline{\cup_{k\in{\mathbb{N}}}Q_{k}}$, where
@@
For $M\geq 2$, by $\tilde{\Omega}$ $M$-tiles $\Omega$ we mean there are $M$ isometries of $\tilde{\Omega}$, $\tilde{\Omega}_{1\leq k\leq M}$, with non-overlapped interior, such that $\cup_{1\leq k\leq M}\tilde{\Omega}_{k}\subset\Omega$ and $|\Omega|=M|\tilde{\Omega}|$.
@@ 1.4 Our method and plan of the paper
Although only a part of discussions in Sections 2 and 3 will be sufficient for the proofs of main results, the extended discussions are of independent interest and may be used for other purposes. For Theorem 1.5 we shall apply the Whitney decomposition of a Lipschitz domain $\Omega$, and using the lower bound for the counting function on cubes (see Lemma 3.2 and Lemma 3.5 below) along with an optimization to deduce a lower bound for the counting function $\mathcal{N}_{\Omega}^{D}(\lambda)$ on $\Omega$. Finally, applying our refinements on product domains (see Theorem 2.3 below) and monotonicity of Dirichlet eigenvalues, together with the lower bound of $\mathcal{N}_{\mathscr{R}\setminus\overline{\Omega}}^{D}(\lambda)$ to conclude the proof. Theorem 1.3 then follows almost immediately from Theorem 1.5.
Finally, let us point it out, our approach depends heavily on the min-max principle (monotonicity of Dirichlet eigenvalues), which is close to the original one by Courant [5] and Courant-Hilbert [6]. However, the way we use the min-max principle is rather different from that in Courant [5] and Courant-Hilbert [6]. By our choose of minimal Admissible rectangles, we do not even have to take care of the cubes which intersect the boundary of $\Omega$. The latter is the main issue for Courant [5] and Courant-Hilbert [6].
@@
In Section 3, we provide the proof for the results on concrete examples of Pólya’s conjecture, Theorem 1.8. Since in higher dimensions, $n\geq 3$, we can replace the rectangle $\mathscr{R}_{n-1}$ by general set satisfying Pólya’s conjecture, we shall split the proof into the case of two dimension, and three and higher dimensions where we prove a more general result, Theorem 3.6. We shall also provide a qualitative result on removing general piecewise smooth set, Theorem 3.1, and shall show some results regarding cubes of Admissible class from some more general class of domains, see Theorem 3.7 below.
@@
For $C>0$ we use $\lfloor C\rfloor$ to denote the integer part of it. For a cube $Q$ we denote by $\ell(Q)$ its side length. For a bounded open domain $\Omega$, we denote by $r_{in}=r_{in}(\Omega)$ the radius of the largest ball contained in $\Omega$. Throughout the paper, most of constants are given explicitly, if not we shall point it out.
@@ 2 Refined eigenvalue estimates on product domains
Let us consider the ball $B(0,\frac{|\Omega|}{\pi}\sqrt{\lambda})$ in the first quadrant in ${\mathbb{R}}^{2}$. For each $1\leq k\leq k_{\lambda}=\lfloor\frac{|\Omega|}{\pi}\sqrt{\lambda}\rfloor$, $(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k^{2})^{1/2}$ equals the area of the rectangle $[k-1,k]\times[0,(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k^{2})^{1/2}]$, which is contained in $B(0,\frac{|\Omega|}{\pi}\sqrt{\lambda})$ in the first quadrant in ${\mathbb{R}}^{2}$. Moreover, for each $k$, there is at least a triangle with vertices at $(k-1,(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-(k-1)^{2})^{1/2})$, $(k-1,(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k^{2})^{1/2})$ and $(k,(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k^{2})^{1/2})$ contained in the ball. Moreover, if $k_{\lambda}\neq\frac{|\Omega|}{\pi}\sqrt{\lambda}$, there is another triangle insider the 1/4 ball with vertices at $(k_{\lambda},(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k_{\lambda}^{2})^{1/2})$, $(k_{\lambda},(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k_{\lambda}^{2})^{1/2})$ and $(\frac{|\Omega|}{\pi}\sqrt{\lambda},0)$. See Figure 2111Picture drawn by Xinyi Jiang..
@@
$\frac{\pi}{|\Omega|}\sum_{1\leq k\leq k_{\lambda}}(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k^{2})^{1/2}<\frac{|\Omega|}{4}\lambda-\frac{\sqrt{\lambda}}{2}+\frac{\sqrt{6}\pi}{9|\Omega|}\sqrt{\lfloor\frac{|\Omega|}{\pi}\sqrt{\lambda}\rfloor+\frac{1}{2}}=\frac{|\Omega|}{4}\lambda\left(1-\frac{2}{3}\frac{1}{|\Omega|\sqrt{\lambda}}\right).$
@@
In view of Laptev’s beautiful theory on product domains [24] (which has been further employed in Laptev-Weidl [25] for Schrödinger operator), the Aizenman-Lieb principle [1], we have the following refined eigenvalue estimates on product domains if one domain satisfies Pólya’s conjecture. Note that any bounded domain $\Omega\subset{\mathbb{R}}$ satisfies Pólya’s conjecture, i.e., $\lambda_{k}(\Omega)\geq\frac{\pi^{2}}{|\Omega|^{2}}k^{2}$ for each $k\in{\mathbb{N}}$.
@@
Then for each $1\leq k\leq k_{0}$, the Pólya conjecture holds on $\Omega$ for $\lambda_{k}(\Omega)$, i.e.,
@@
Let $1\leq k\leq k_{0}$. If $\lambda_{k}(\Omega)<\lambda_{0}$, then (2) gives that for small enough $\epsilon>0$ such that $\lambda_{k}(\Omega)+\epsilon<\lambda_{0}$
@@
Let us consider $\Omega_{1}\subset{\mathbb{R}}^{2}$, with $|\Omega_{1}|=100$, $\Omega_{2}\subset{\mathbb{R}}$ with $|\Omega_{2}|=1$. Suppose that Pólya’s conjecture holds on $\Omega_{1}$ for the Dirichlet eigenvalues. In this case, $C_{1}(2)=3\pi/16$,

### v2

@@
where $\Lambda(\epsilon,\Omega)$ is given explicitly. This estimate is based on a uniform estimate on the remainder of the Weyl law with explicit constants so that it reduces the $\epsilon$-loss version of Pólya’s conjecture to a computational problem. Our arguments in deriving such uniform estimates yield also, in all dimensions $n\geq 2$, classes of domains that may even have rather irregular shapes or boundaries but satisfy Pólya’s conjecture.
@@ 1.1 Weyl’s law and Pólya’s conjecture
The Weyl law is also important in the longstanding question “can one hear about the shape of a cavity”, see Lord Rayleigh [30] and Kac [23]. Moreover, the Weyl law maybe related to the prime number theorem, via Connes’ trace formula, see Connes [4], Fathizadeh-Khalkhali [13] and the beamer by Khalkhali [24].
@@
Subsequent to Weyl’s seminal work, Courant [5], Courant-Hilbert [6], Hörmander [19, 20], Seeley [35, 36] and many others (see Ivrii [22] for comprehensive description on related history) made contributions to improve the estimate on the remainder.
@@
However, the above mentioned results for the remainder $R_{\Omega}(\lambda)$ state that certain asymptotic estimates valid only for large $\lambda$’s without quantifying the sizes of such $\lambda$’s, and consequently they do not give quantitative estimate on eigenvalues. [^1] After the first version of the present article was posted on arXiv, R. Frank and S. Larson has kindly informed us recent works [9] and [12], in which several uniform estimates for the remainder were already established, see the added remark at the end of the introduction.
It is rather natural to ask the following question.
@@
In 1954, Pólya [32] conjectured that, it should hold on arbitrary bounded open domain that
@@
The Pólya conjecture has been wide open since then. Other than tiling domains, Laptev [26] verified Pólya’s conjecture on product domains $\Omega_{1}\times\Omega_{2}\subset{\mathbb{R}}^{n_{1}+n_{2}}$, $n_{1}\geq 2$, $n_{2}\geq 1$, provided Pólya’s conjecture holds on $\Omega_{1}$. Filonov et al. [7] and [8] verified, via number theoretic arguments, Pólya’s conjecture on balls for all $n\geq 2$, and on annuli for $n=2$. For general domains $\Omega$, it was known that Pólya’s conjecture holds for the first two Dirichlet eigenvalues see [18] for the Rayleigh-Faber-Krahn inequality ($\lambda_{1}(\Omega)$) and the Krahn-Szegö inequality ($\lambda_{2}(\Omega)$). See also [18] for the first nontrivial Neumann eigenvalue and recent exciting developments [3, 15] for the second nontrivial Neumann eigenvalue.
@@
Our main purpose of this paper is to reduce the $\epsilon$-loss version of Pólya’s conjecture to a computational problem, and to give a solution to Question 1.1 with explicit constants, on Lipschitz domains. Moreover, we shall also provide an approach to the Pólya conjecture that might be applied to general smooth domains, and provide in all dimensions, $n\geq 2$, a class of domains that satisfy Pólya’s conjecture.
1.2 Pólya’s conjecture up to $\epsilon$-loss and quantitative remainder estimate
Let us start from a definition for Pólya’s conjecture up to an $\epsilon$-loss.
@@ 1.2 Pólya’s conjecture up to $\epsilon$-loss and quantitative remainder estimate
Noting that $2\leq\left(\frac{n+2}{n}\right)^{n/2}\to e$ as $2\leq n\to\infty$, the $\epsilon$-loss version of Pólya’s conjecture improves substantially (1.5) (and its improvements on lower order terms, see [2, 14, 25, 31]).
@@
We have the following Pólya’s conjecture up to $\epsilon$-loss.
@@
Moreover, if $\Omega$ is convex, then $\Lambda(\epsilon,\Omega)$ can be taken smaller as
(1.8) $\frac{2n\mathrm{diam}(\Omega)^{n-1}}{|\Omega|\sqrt{\Lambda(\epsilon,\Omega)}}\left(\frac{\pi}{2}+5n^{2}\pi\log_{2}\frac{10\mathrm{width}(\Omega)\sqrt{\Lambda(\epsilon,\Omega)}}{\pi}\right)=\epsilon.$
Obviously, for a given $\epsilon\in(0,1)$, and a concrete Lipschitz (resp. convex) domain $\Omega$, $\Lambda(\epsilon,\Omega)$ is uniquely determined by the equation (1.7) (resp. (1.8) for convex domains). It can be seen from the above estimate that $\Omega$ needs only to be a bounded set of finite perimeter. Also, to show the $\epsilon$-loss version of Pólya’s conjecture, Theorem 1.3 reduces the problem to a computational problem for eigenvalues $\lambda_{k}(\Omega)<\Lambda(\epsilon,\Omega)$.
@@
We have the following uniform estimate on Lipschitz domains. In what follows, for a bounded open domain $\Omega$, we denote by $r_{in}=r_{in}(\Omega)$ the radius of the largest ball contained in $\Omega$.
@@
(1.10) $k$ $\geq\frac{\omega(n)|\Omega|}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2}-\frac{C_{Lip}(\Omega)\omega(n)|\partial\Omega|}{(2\pi)^{n}}\lambda_{k}(\Omega)^{\frac{n-1}{2}}\left(\frac{\pi r_{in}(\Omega)}{|I_{min}|}+5n^{2}\pi\log_{2}\frac{10|I_{min}|\sqrt{\lambda_{k}(\Omega)}}{\pi}\right).$
@@
Obviously, both quantities $\frac{r_{in}(\mathscr{R}\setminus\overline{\Omega})}{|I_{min}|}$ and $\frac{r_{in}(\Omega)}{|I_{min}|}$ are not larger than $1/2$, and therefore can be replaced by $1/2$. Moreover, note that if $|\partial(\mathscr{R}\setminus\overline{\Omega})|$ is small compared to $|\mathscr{R}_{n-1}|$, we then have that Pólya’s conjecture holds for the first some eigenvalues. Our quantitative estimate is consistent to the asymptotic estimates established by Courant [5] and Courant-Hilbert [6].
For convex sets, we have a clean and better estimate as following.
Corollary 1.6.
Let $\Omega\subset{\mathbb{R}}^{n}$ be a bounded convex domain, $n\geq 2$. Let $\mathscr{R}=\mathscr{R}_{n-1}\times I_{min}$ be the minimal Admissible rectangle for $\Omega$. It holds for all $k\in{\mathbb{N}}$ that
$k$ $\leq\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2}$
$\quad+\frac{\omega(n)}{(2\pi)^{n}}\lambda_{k}(\Omega)^{\frac{n-1}{2}}\left(|\partial\mathscr{R}\setminus\partial\Omega|\left(\frac{\pi r_{in}(\mathscr{R}\setminus{\overline{\Omega}})}{|I_{min}|}+5n^{2}\pi\log_{2}\frac{10|I_{min}|\sqrt{\lambda_{k}(\Omega)}}{\pi}\right)-C_{1}(n-1)|\mathscr{R}_{n-1}|\right),$
and
$k$ $\geq\frac{\omega(n)|\Omega|}{(2\pi)^{n}}\lambda_{k}(\Omega)^{n/2}-\frac{\omega(n)|\partial\Omega|}{(2\pi)^{n}}\lambda_{k}(\Omega)^{\frac{n-1}{2}}\left(\frac{\pi r_{in}(\Omega)}{|I_{min}|}+5n^{2}\pi\log_{2}\frac{10|I_{min}|\sqrt{\lambda_{k}(\Omega)}}{\pi}\right).$
@@
1.3 An approach to Pólya’s conjecture and concrete examples
Using the strategy developed above, we can provide in all dimensions $n\geq 2$, a class of domains which may have rather irregular shape or boundary but satisfy Pólya’s conjecture. Using Seeley’s asymptotic estimates [35, 36] (see also Hörmander [20]), we first provide an approach to Pólya’s conjecture that might be applied to general smooth domains.
Theorem 1.7.
Suppose $\mathscr{R}=\mathscr{R}_{n-1}\times I$ is a rectangle in ${{{\mathbb{R}}}^{n}}$, $n\geq 2$. Assume that $\Omega$ is a piecewise smooth domain contained in $\mathscr{R}$. There exists $C_{P}=C(\Omega,n)$ such that if $|\mathscr{R}_{n-1}|\geq C_{P}$ then the Pólya conjecture holds on $\mathscr{R}\setminus\overline{\Omega}$.
@@ 1.3 An approach to Pólya’s conjecture and concrete examples
We have a class of domains that satisfy Pólya’s conjecture in all dimensions.
@@
Suppose $\mathscr{R}=\mathscr{R}_{n-1}\times I$ is a rectangle in ${{{\mathbb{R}}}^{n}}$, $n\geq 2$. Let $\{Q_{k}\}_{k\in{\mathbb{N}}}$ be of Admissible class in ${{{\mathbb{R}}}^{n}}$, that is contained in $\mathscr{R}$. If $|\mathscr{R}_{n-1}|\geq C_{2}(n-1)\mathscr{S}_{Q}$, then the Pólya conjecture holds on $\mathscr{R}\setminus\overline{\cup_{k\in{\mathbb{N}}}Q_{k}}$, where
@@
For $M\geq 2$, by $\tilde{\Omega}$ $M$-tiles $\Omega$ we mean there are $M$ isometries of $\tilde{\Omega}$, $\{\tilde{\Omega}_{k}\}_{1\leq k\leq M}$, with non-overlapped interior, such that $\cup_{1\leq k\leq M}\tilde{\Omega}_{k}\subset\Omega$ and $|\Omega|=M|\tilde{\Omega}|$.
@@ 1.4 Our method and plan of the paper
Although only a part of discussions in Sections 2 and 3 will be sufficient for the proofs of Theorem 1.3 and Theorem 1.5, the extended discussions are of independent interest and they have other uses. For Theorem 1.5 we shall apply the Whitney decomposition of a Lipschitz domain $\Omega$, and using the lower bound for the counting function on cubes (see Lemma 3.3 and Lemma 3.6 below) along with an optimization to deduce a lower bound for the counting function $\mathcal{N}_{\Omega}^{D}(\lambda)$ on $\Omega$. Finally, applying our refinements on product domains (see Theorem 2.3 below) and monotonicity of Dirichlet eigenvalues, together with the lower bound of $\mathcal{N}_{\mathscr{R}\setminus\overline{\Omega}}^{D}(\lambda)$ to conclude the proof. Theorem 1.3 then follows almost immediately from Theorem 1.5.
Finally, let us point it out, our approach depends on a modified min-max argument, which is related to but different from the original ones by Courant [5] and Courant-Hilbert [6]. By our choose of minimal Admissible rectangles, we do not have to take care of the term in Courant [5] and Courant-Hilbert [6] for cubes which intersect the boundary of $\Omega$. The latter is a difficult issue for Courant [5] and Courant-Hilbert [6].
@@
In Section 3, we provide the proofs for Theorem 1.7 and the results on concrete examples of Pólya’s conjecture, Theorem 1.9. Since in higher dimensions, $n\geq 3$, we can replace the rectangle $\mathscr{R}_{n-1}$ by general set satisfying Pólya’s conjecture, we shall split the proof into the case of two dimension, and three and higher dimensions where we prove a more general result, Theorem 3.7. We shall also show some results regarding cubes of Admissible class from some more general class of domains, see Theorem 3.8 below.
@@
For $C>0$ we use $\lfloor C\rfloor$ to denote the integer part of it. For $a\in{\mathbb{R}}$ we denote by $a_{+}$ the quantity $\max\{0,a\}$. For a cube $Q$ we denote by $\ell(Q)$ its side length. For a bounded open domain $\Omega$, we denote by $r_{in}=r_{in}(\Omega)$ the radius of the largest ball contained in $\Omega$. Throughout the paper, most of constants are given explicitly, if not we shall point it out.
Added in remark. After the first version of this article was posted on arXiv, we are kindly informed by Frank and Larson, that a similar version to our uniform remainder estimate for Weyl’s law (Theorem 1.5) already appears in [9, Theorem 3.14] and [12, Proposition 3.6], which both show that
(1.13) $\left|\mathcal{N}^{D}_{\Omega}(\lambda)-\frac{|\Omega|\omega(n)}{(2\pi)^{n}}\lambda^{n/2}\right|$ $\leq C(n)\Theta_{\Omega}\lambda^{\frac{n-1}{2}}\left(1+\log_{+}(r_{in}(\Omega)\sqrt{\lambda})\right),$
where $C(n)$ is an implicit constant depending only on the dimension $n$,
$\Theta_{\Omega}:=\sup_{\delta>0}\delta^{-1}|\left\{x\in\Omega:\,{\mathop{\mathrm{missing}}{\,dist\,}}(x,\partial\Omega)<\delta\right\}|<\infty,$
$\Theta_{\Omega}$ plays the same role of $C_{Lip}(\Omega)|\partial\Omega|$ in our estimate. In particular, it appears from the proof of [9, Theorem 3.14] that the above constant $C(n)$ can also be calculated. The proof of [12, Proposition 3.6] uses Tauberian theory, while the proof of [9, Theorem 3.14] uses Courant’s method of Dirichlet-Neumann bracketing and some refined estimates for cubes intersecting the boundary.
Our proof for estimates of the reminder of Weyl’s law (Theorem 1.5) uses the minimal Admissible rectangle (see Theorem 2.3 below for refined eigenvalue estimates on product domains), a subtracting argument that allows us to avoid the cubes intersecting the boundary, and then lower bound for the counting function of Dirichlet eigenvalues. Moreover, our method has an advantage that gives many “irregular shaped domains” that verify the Pólya conjecture (see Theorem 1.9, Remark 1.11 and Theorem 3.7 below).
@@ 2 Refined eigenvalue estimates on product domains
Let us consider the ball $B(0,\frac{|\Omega|}{\pi}\sqrt{\lambda})$ in the first quadrant in ${\mathbb{R}}^{2}$. For each $1\leq k\leq k_{\lambda}=\lfloor\frac{|\Omega|}{\pi}\sqrt{\lambda}\rfloor$, $(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k^{2})^{1/2}$ equals the area of the rectangle $[k-1,k]\times[0,(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k^{2})^{1/2}]$, which is contained in $B(0,\frac{|\Omega|}{\pi}\sqrt{\lambda})$ in the first quadrant in ${\mathbb{R}}^{2}$. Moreover, for each $k$, there is at least a triangle with vertices at $(k-1,(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-(k-1)^{2})^{1/2})$, $(k-1,(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k^{2})^{1/2})$ and $(k,(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k^{2})^{1/2})$ contained in the ball. Moreover, if $k_{\lambda}\neq\frac{|\Omega|}{\pi}\sqrt{\lambda}$, there is another triangle insider the 1/4 ball with vertices at $(k_{\lambda},(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k_{\lambda}^{2})^{1/2})$, $(k_{\lambda},(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k_{\lambda}^{2})^{1/2})$ and $(\frac{|\Omega|}{\pi}\sqrt{\lambda},0)$. See Figure $2$Picture drawn by Xinyi Jiang..
@@
$\frac{\pi}{|\Omega|}\sum_{1\leq k\leq k_{\lambda}}(\frac{|\Omega|^{2}}{\pi^{2}}\lambda-k^{2})^{1/2}<\frac{|\Omega|}{4}\lambda-\frac{\sqrt{\lambda}}{2}+\frac{\sqrt{6}\pi}{9|\Omega|}\sqrt{\lfloor\frac{|\Omega|}{\pi}\sqrt{\lambda}\rfloor+\frac{1}{2}}\leq\frac{|\Omega|}{4}\lambda\left(1-\frac{2}{3}\frac{1}{|\Omega|\sqrt{\lambda}}\right).$
@@
In view of Laptev’s beautiful theory on product domains [26] (which has been further employed in Laptev-Weidl [27] for Schrödinger operator), the Aizenman-Lieb principle [1], we have the following refined eigenvalue estimates on product domains if one domain satisfies Pólya’s conjecture. We note that Larson [28] earlier had obtained similar improvements for higher dimensions ($n_{1}\geq 3$ below) for more general products but with less explicit constants. Note that any bounded domain $\Omega\subset{\mathbb{R}}$ satisfies Pólya’s conjecture, i.e., $\lambda_{k}(\Omega)\geq\frac{\pi^{2}}{|\Omega|^{2}}k^{2}$ for each $k\in{\mathbb{N}}$.
@@
Then for each $0\leq k\leq k_{0}$, the Pólya conjecture holds on $\Omega$ for $\lambda_{k}(\Omega)$, i.e.,
@@
Let $0\leq k\leq k_{0}$. If $\lambda_{k}(\Omega)<\lambda_{0}$, then (2) gives that for small enough $\epsilon>0$ such that $\lambda_{k}(\Omega)+\epsilon<\lambda_{0}$
@@
Obviously the above corollary only makes sense when $k_{0}\geq 3$. Let us consider $\Omega_{1}\subset{\mathbb{R}}^{2}$, with $|\Omega_{1}|=100$, $\Omega_{2}\subset{\mathbb{R}}$ with $|\Omega_{2}|=1$. Suppose that Pólya’s conjecture holds on $\Omega_{1}$ for the Dirichlet eigenvalues. In this case, $C_{1}(2)=3\pi/16$,